OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. 8x + y = 10 6x + 5y = -1

OpenStudy (anonymous):

You need to substitute for X and Y.

OpenStudy (anonymous):

how

OpenStudy (anonymous):

@TrashMan914 I am not great at substitution would you mind helping her and I?

OpenStudy (anonymous):

yes. please help

OpenStudy (anonymous):

Calculated something wrong. Going to redo this problem.

OpenStudy (anonymous):

okayy thank you

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

hes working it out I don't know this to well

OpenStudy (anikhalder):

Multiply eqn 1 with 5 and then subtract eqn 2 from it. 40x + 5y=50 - 6x + 5y=-1 Which results to 34x=51 => x=3/2 Now substitute the value of x in any eq. say 2.. 6*(3/2) +5y=-1 =>9+5y=-1 =>5y=-10 =>y=-2 So, x=3/2 and y=-2 We may easily verify and check the answers

OpenStudy (anonymous):

ohhhhhh i understand now

OpenStudy (anonymous):

wait no.

OpenStudy (anikhalder):

then?

OpenStudy (anonymous):

idk.

OpenStudy (anonymous):

im so lostt

OpenStudy (anikhalder):

Where's the difficulty you are facing? Feel free to ask